# MoCHA: Advanced Vision-Language Reasoning with MoE Connector and Hierarchical Group Attention

### v2

@@ Introduction
Vision Large Language Models (VLLMs) (Liu et al. 2024b; Wang et al. 2024b; Bai et al. 2025) have shown promising results in multimodal reasoning tasks by integrating visual encoders with Large Language Models (LLM) (OpenAI 2023, 2024; Touvron et al. 2023). However, the large model size and extensive training data present significant computational challenges. Additionally, excessive early reliance on language in VLLMs may act as a shortcut (Geirhos et al. 2020; Yuksekgonul et al. 2023) that results in suboptimal learning of effective visual representations. Therefore, exploring a VLLM that balances efficient visual processing and performance is a crucial topic.
@@
Previous works have focused on strengthening vision encoders by utilizing extensive and higher-quality visual instruction data (Zheng et al. 2023), scaling up model size (Chen et al. 2024), or decomposing images into low-resolution patches (Shi et al. 2024). While effective, these approaches are inherently limited by high computational costs and the loss of important details, such as small objects, which in turn can result in hallucinations. Furthermore, Cambrian-1 (Tong et al. 2024), designed with a vision centric approach, introduces a novel paradigm for vision-language fusion through its combination of multiple vision encoders and spatial vision aggregator (SVA), but faces difficulties due to the lack of dynamic feature fusion and shortcomings in fine-grained perception. Specifically, visual features are inherently high-dimensional and uncertain, encompassing objects, scenes, attributes, and spatial relationships. A single vision encoder, or a limited set thereof, is insufficient to comprehensively and adaptively capture these diverse aspects. Sparse Mixture of Experts (MoE) (Fedus, Zoph, and Shazeer 2022) has emerged as an efficient solution that dynamically activates only a subset of expert networks for each input to improve scalability and specialization. Although MoE has been widely adopted in state-of-the-art open-source LLMs, its potential remains largely underexplored in the design of connectors and the broader architecture of VLLMs.
@@
To address the above limitations of existing vision encoders and enhance fine-grained visual understanding, we propose MoCHA, a novel framework integrates multiple vision backbones under a sparse Mixture of Experts Connector (MoECs) module and a Hierarchical Group Attention (HGA) component for vision-extensive reasoning tasks. Specifically, MoCHA consists of multiple vision encoders connected by a mixture of experts module and hierarchical group attention, as illustrated in Figure 1. First, we leverage four distinct yet complementary vision backbones, including CLIP (Radford et al. 2021), SigLIP (Zhai et al. 2023), DINOv2 (Oquab et al. 2024), and ConvNeXt (Liu et al. 2022), to extract diverse and robust image features. However, the efficient integration of these heterogeneous visual signals into a unified vision-language framework remains a challenge. To address this, we propose the mixture of experts connectors (MoECs) module, in which each connector is implemented as a Top-K sparsely-gated MoE for dynamic expert selection across visual dimensions. This facilitates efficient cross-modal interaction and reduces training costs. Sequential concatenation of features from multiple vision encoders along the token dimension often leads to redundancy and feature overlap. We further design hierarchical group attention (HGA), which fuses features through intra- and inter-group attention. By adaptive gating mechanism, we dynamically balance the aggregated and original features to produce the image representation without additional parameters. Our MoCHA is fully trained on open-source datasets and demonstrates superior performance compared to other VLLMs of similar size. In detail, we implement MoCHA with a two-stage training process on Phi2-2.7B and Vicuna-7B LLMs, achieving higher scores than existing open-weight VLLMs on mainstream vision-language benchmarks. Notably, our 3B-sized MoCHA shows a 3.25% reduction in hallucination on POPE and delivers a 153-point improvement on general visual tasks on MME, outperforming the larger CuMo-7B model (Li et al. 2024a). This demonstrates the efficacy of our MoCHA in mitigating VLM hallucinations and enhancing general perceptual abilities. We further conduct ablation studies to investigate the contribution of each component to overall performance.
@@
The advent of pre-trained Vision Transformers (ViT) (Dosovitskiy et al. 2021) has significantly propelled the advancement of computer vision. As the original CLIP adopted by conventional visual instruction tuning approaches is trained on noisy image-text pairs, it exhibits specific visual shortcomings, and thus stronger backbones have been introduced to LLMs. For example, SigLIP introduced pairwise sigmoid loss during training, enabling the vision encoder to demonstrate more advanced visual perception capabilities. Some concurrent works have leveraged external assistance from vision-only, self-supervised models such as DINOv2, MoCo-v3 (He et al. 2022), and other advanced frameworks  (Caron et al. 2021; Chen et al. 2024), as well as domain-specific expert models. Additionally, ConvNeXt contributes to this progress by incorporating a purely convolutional backbone that achieves Transformer-level performance with improved computational efficiency. Collectively, these approaches foster robust multimodal integration and significantly enhance visual reasoning capabilities in contemporary VLLM frameworks by leveraging complementary vision encoders to enrich visual representations and improve fine-grained perception.
@@
Mixture of Experts (MoE) proposes a group of expert networks to handle specific tasks, with a gating network selecting the appropriate experts. Subsequent research on language MoE models has further expanded large MoE-based language models, improving expert stability and load balancing. ST-MoE (Zoph et al. 2022) employs load-balancing loss and router-z loss to ensure an even distribution of experts. The recently popular DeepSeek-V3 adopts the DeepSeekMoE (Dai et al. 2024) architecture, leveraging shared experts to capture common knowledge and reduce redundancy in routed experts. MoE’s success has also extended to the vision domain. LIMoE (Mustafa et al. 2022) replaces dense MLP layers with MoE layers in CLIP, improving zero-shot image classification. AdaMV-MoE (Liu and Luo 2024) introduces an adaptive MoE framework for multi-task learning. Furthermore, CuMo (Li et al. 2024a) integrates Co-upcycled Top-K sparse-gating MoE blocks into the visual encoder, MLP connector, and language model, enhancing VLLM inference performance.
@@
To address a critical gap where current VLLMs still struggle—solving vision-extensive multimodal tasks—we present our framework, MoCHA, in this section. Building upon the mainstream LLaVA architecture, MoCHA moves beyond the original CLIP encoder by integrating a diverse set of vision encoders, spanning different architectures and pre-training objectives. This enables the model to effectively capture and process visual information across varying levels of granularity. As illustrated in Figure 1, our framework consists of multiple vision encoders (i.e., OpenAICLIP ViT-L/14@336, SigLIP ViT-L/16@384, OpenCLIP ConvNeXt-XXL@1024 and DINOv2 ViT-L/14@336) connected by a mixture of experts module and hierarchical group attention.
@@
Although language-supervised models outperform self-supervised and other models across all benchmark categories, Cambrian-1 (Tong et al. 2024) highlights that well-trained self-supervised models like DINOv2 are capable of achieving competitive performance in vision-centric tasks. In particular, high-resolution models excel in chart-related and other vision-centric benchmarks while also demonstrating robust performance across both general VQA and knowledge-based VQA. While ViTs remain the dominant architecture, ConvNet-based models such as OpenCLIP ConvNeXt are well-suited for high-resolution image processing (Vishniakov, Shen, and Liu 2024) delivering outstanding results on OCR, chart interpretation, and other vision-centric benchmarks.
@@
Previous mainstream approaches (Shazeer et al. 2017) replace dense MLP blocks with sparsely-gated mixture of experts blocks. As an example, consider the SigLIP vision encoder embedding. Given input $e_{s}\in\mathbb{R}^{s\times D_{s}}$ and an MLP block, the hidden representation is computed as:
@@
While channel concatenation (Lin et al. 2023) achieves performance comparable to sequence append (Liu et al. 2024c; Kar et al. 2024; Fan et al. 2024), it requires strict spatial alignment of encoder outputs, rendering it less compatible with encoders of differing resolutions and architectures. Sequence append along the token dimension obviates token count alignment and removes the requirement for interpolation, resampling, or padding. As a result, encoders with different resolutions, patch sizes, or feature map dimensions can contribute arbitrary numbers of tokens. In this study, we adopt sequence append, concatenating features from different encoders directly along the token dimension:
@@
To maintain a load balance among experts in each MoEC block, we adopt auxiliary losses based on the language modeling cross-entropy loss. This auxiliary loss consists of a load balancing loss (Wang et al. 2024a) and a router z-loss (Zoph et al. 2022). Hence, the total loss is:
@@
 Qwen-VL-Chat  (Bai et al. 2023)  Qwen-7B    57.5 68.2 61.5 - - 1487.5 60.6 -
 InstructBLIP  (Dai et al. 2023)  Vicuna-7B    49.2 60.5 50.1 26.2 - - 36.0 -
 LLaVA-LLaMA3 (Contributors 2023)  LLaMA3-8B-IT    62.6 72.9 59.0 - 86.4 1469 72.3 -
 Mini-Gemini (Li et al. 2024b)  Vicuna-7B    - 65.2 - 40.8 - 1523 69.3 31.4
 SPHINX-Intern2 (Liu et al. 2024a)  InternLM2-7B    56.2 70.4 58.1 36.5 86.9 1260.4 57.9 35.5
 LLaVA-NeXT (Liu et al. 2024b)  Vicuna-7B    64.2 70.1 64.9 43.9 86.5 1519 67.4 34.6
 LLaVA-NeXT (Liu et al. 2024b)  Mistral-7B    64.8 72.8 65.7 47.3 86.7 1498 68.7 37.7
 LLaVA-v1.5 (Liu et al. 2023)  Vicuna-7B    62.0 66.8 58.2 30.5 85.9 1510.7 64.3 -
 VILA (Lin et al. 2024b)  Vicuna-7B    62.3 68.2 64.4 34.9 85.5 1533 68.9 -
 CuMo (Li et al. 2024a)  Mistral-7B    64.9 73.9 67.0 51.0 86.7 1548.6 73.0 35.1
 InstructBLIP (Dai et al. 2023)  Vicuna-13B    49.5 63.1 50.7 25.6 78.9 1212.8 - -
 LLaVA-v1.5 (Liu et al. 2023)  Vicuna-13B    63.3 71.6 61.3 35.4 85.9 1531.3 67.7 27.6
 Mini-Gemini (Li et al. 2024b)  Vicuna-13B    - 65.9 - 46.0 - 1565 68.5 37.0
 InternVL-Chat (Chen et al. 2024)  Vicuna-13B    66.6 - 61.5 - 87.6 1586.4 - -
 LLaMA-VID (Li, Wang, and Jia 2024)  Vicuna-13B    65.0 70.0 - - 86.0 1542.3 66.6 -
 SPHINX-Plus (Liu et al. 2024a)  LLaMA2-13B    - 74.2 65.7 47.9 89.1 1457.7 71.0 36.8
 MobileVLM (Chu et al. 2023)  MobileLLaMA-2.7B    59.0 61.0 47.5 - 84.9 1288.9 59.6 -
 TinyGPT-V (Yuan et al. 2024)  Phi2-2.7B    33.6 - - - - - - -
 LLaVA-Phi (Zhu et al. 2024)  Phi2-2.7B    - 68.4 48.6 28.9 85.0 1335.1 59.8 -
 MoE-LLaVA-2.7B×4-Top2 (Lin et al. 2024a)  Phi2-2.7B    61.4 68.5 51.4 34.3 86.3 1423 65.2 -
@@
We focus on academic VQA datasets, including GQA (Hudson and Manning 2019), Science-QA (Lu et al. 2022), and TextVQA (Singh et al. 2019), as well as instruction-following VLLM benchmarks, such as POPE (Li et al. 2023), MME (Fu et al. 2024), MMBench (Liu et al. 2024d), and MM-Vet (Yu et al. 2024). Furthermore, we evaluate the challenging MathVista (Lu et al. 2024) dataset to assess VLLM visual reasoning capabilities.
@@
As we scale up the LLM backbone, MoCHA (Vicuna-7B) demonstrates stronger visual reasoning capabilities than MoCHA (Phi-2 2.7B). Notably, MoCHA (Vicuna-7B) also surpasses models with even larger LLMs across multiple visual benchmarks. For instance, compared to LLaVA-v1.5 (Vicuna-13B), it achieves improvements of 2.99%, 4.42%, and 4.08% on SQA IMG, TextVQA, and POPE, respectively. Moreover, in contrast to CuMo—which applies a global mixture of experts (MoE) across the LLM, vision encoder, and connector—MoCHA (Vicuna-7B) outperforms CuMo (Mistral-7B) on several key vision benchmarks. These results provide strong evidence that, compared to the conventional single-vision-encoder paradigm, MoCHA’s dynamic MoE design and customized attention mechanisms offer a more effective approach for multimodal learning. We present more results regarding other visual capabilities such as multi-object recognition in Appendix D.
@@
To validate the efficiency use of our MoECs, we further present training GFLOPs, parameters, and average inference time of models in Table 2. We make sure that all numbers are obtained under identical configurations to ensure the validity of the comparison. From the table, MoCHA achieves an inference speed of 0.57s with only 4.97B parameters and 12014.64 GFLOPs. This not only significantly reduces both the parameter size and computational cost, but also outperforms LLaVA-v1.5 (Vicuna-13B) and InternVL-Chat (Vicuna-13B) in inference efficiency. Furthermore, it approaches the inference speed of MoE-LLaVA, despite using even fewer parameters. These notable advantages in both computational and parameter efficiency highlight the strong potential of scaling up MoCHA as a compelling alternative to existing vision encoder paradigms in multimodal learning.
@@
Ablation Studies
Effect of Channel v.s. Sequential Concatenation.
@@
SigLIP+CLIP+ConvNeXt    59.1 66.58 87.06 60.83
@@
To evaluate MoEC’s capability in efficiently integrating heterogeneous and multi-dimensional features, we initiate this ablation study by replacing each MLP connector with an upcycled MoEC, as shown in Table 4. The results in Table 3 indicate that with the plain MLP as the connector, the model performance with a single SigLIP remains weak. While the addition of ConvNeXt further improves performance, the full integration of CLIP yields the largest gain among all settings. On the other hand, in Table 4, the combination of SigLIP, DINOv2, ConvNeXt, and CLIP under our MoEC results in notable performance improvements of 1.49%, 1.91%, 0.54%, and 1.52% on the GQA, SQA IMG, POPE, and MMBench benchmarks, respectively, compared to results using vanilla MLP in Table 3. These findings show that MoECs effectively handle diverse visual features—including object recognition, scene understanding, attribute identification, and spatial reasoning. Even as the number of vision encoders grows, MoECs incur minimal parameter overhead while keeping inference time and computational cost on par with a standard MLP. Further quantitative results are provided in Appendix C.
Method (MoEC + Attention)    GQA

### v3

@@ Introduction
Vision Large Language Models (VLLMs) (liu2024llavanext; wang2024qwen2vlenhancingvisionlanguagemodels; bai2025qwen25vltechnicalreport) have shown promising results in multimodal reasoning tasks by integrating visual encoders with Large Language Models (LLM) (gpt; openai2024gpt4o; touvron2023llama). However, the large model size and extensive training data present significant computational challenges. Additionally, excessive early reliance on language in VLLMs may act as a shortcut (geirhos2020shortcut; yuksekgonul2023visionlanguagemodelsbehavelike) that results in suboptimal learning of effective visual representations. Therefore, exploring a VLLM that balances efficient visual processing and performance is a crucial topic.
@@
Previous works have focused on strengthening vision encoders by utilizing extensive and higher-quality visual instruction data (zheng2023judgingllmasajudgemtbenchchatbot), scaling up model size (chen2024internvl), or decomposing images into low-resolution patches (shi2024needlargervisionmodels). While effective, these approaches are inherently limited by high computational costs and the loss of important details, such as small objects, which in turn can result in hallucinations. Furthermore, Cambrian-1 (tong2024cambrian1fullyopenvisioncentric), designed with a vision centric approach, introduces a novel paradigm for vision-language fusion through its combination of multiple vision encoders and spatial vision aggregator (SVA), but faces difficulties due to the lack of dynamic feature fusion and shortcomings in fine-grained perception. Specifically, visual features are inherently high-dimensional and uncertain, encompassing objects, scenes, attributes, and spatial relationships. A single vision encoder, or a limited set thereof, is insufficient to comprehensively and adaptively capture these diverse aspects. Sparse Mixture of Experts (MoE) (fedus2022switch) has emerged as an efficient solution that dynamically activates only a subset of expert networks for each input to improve scalability and specialization. Although MoE has been widely adopted in state-of-the-art open-source LLMs, its potential remains largely underexplored in the design of connectors and the broader architecture of VLLMs.
@@
To address the above limitations of existing vision encoders and enhance fine-grained visual understanding, we propose MoCHA, a novel framework integrates multiple vision backbones under a sparse Mixture of Experts Connector (MoECs) module and a Hierarchical Group Attention (HGA) component for vision-extensive reasoning tasks. Specifically, MoCHA consists of multiple vision encoders connected by a mixture of experts module and hierarchical group attention, as illustrated in Figure 1. First, we leverage four distinct yet complementary vision backbones, including CLIP (radford2021learningtransferablevisualmodels), SigLIP (zhai2023sigmoidlosslanguageimage), DINOv2 (oquab2024dinov2learningrobustvisual), and ConvNeXt (liu2022convnet2020s), to extract diverse and robust image features. However, the efficient integration of these heterogeneous visual signals into a unified vision-language framework remains a challenge. To address this, we propose the mixture of experts connectors (MoECs) module, in which each connector is implemented as a Top-K sparsely-gated MoE for dynamic expert selection across visual dimensions. This facilitates efficient cross-modal interaction and reduces training costs. Sequential concatenation of features from multiple vision encoders along the token dimension often leads to redundancy and feature overlap. We further design hierarchical group attention (HGA), which fuses features through intra- and inter-group attention. By adaptive gating mechanism, we dynamically balance the aggregated and original features to produce the image representation without additional parameters. Our MoCHA is fully trained on open-source datasets and demonstrates superior performance compared to other VLLMs of similar size. In detail, we implement MoCHA with a two-stage training process on Phi2-2.7B and Vicuna-7B-v1.5 LLMs, achieving higher scores than existing open-weight VLLMs on mainstream vision-language benchmarks. Notably, our 3B-sized MoCHA shows a 3.25% reduction in hallucination on POPE and delivers a 153-point improvement on general visual tasks on MME, outperforming the larger CuMo-7B model (li2024cumoscalingmultimodalllm). This demonstrates the efficacy of our MoCHA in mitigating VLM hallucinations and enhancing general perceptual abilities. We further conduct ablation studies to investigate the contribution of each component to overall performance.
@@
The advent of pre-trained Vision Transformers (ViT) (dosovitskiy2021imageworth16x16words) has significantly propelled the advancement of computer vision. As the original CLIP adopted by conventional visual instruction tuning approaches is trained on noisy image-text pairs, it exhibits specific visual shortcomings, and thus stronger backbones have been introduced to LLMs. For example, SigLIP introduced pairwise sigmoid loss during training, enabling the vision encoder to demonstrate more advanced visual perception capabilities. Some concurrent works have leveraged external assistance from vision-only, self-supervised models such as DINOv2, MoCo-v3 (he2021maskedautoencodersscalablevision), and other advanced frameworks  (caron2021emergingpropertiesselfsupervisedvision; chen2024internvl), as well as domain-specific expert models. Additionally, ConvNeXt contributes to this progress by incorporating a purely convolutional backbone that achieves Transformer-level performance with improved computational efficiency. Collectively, these approaches foster robust multimodal integration and significantly enhance visual reasoning capabilities in contemporary VLLM frameworks by leveraging complementary vision encoders to enrich visual representations and improve fine-grained perception.
@@
Mixture of Experts (MoE) proposes a group of expert networks to handle specific tasks, with a gating network selecting the appropriate experts. Subsequent research on language MoE models has further expanded large MoE-based language models, improving expert stability and load balancing. ST-MoE (zoph2022stmoedesigningstabletransferable) employs load-balancing loss and router-z loss to ensure an even distribution of experts. The recently popular DeepSeek-V3 adopts the DeepSeekMoE (dai2024deepseekmoeultimateexpertspecialization) architecture, leveraging shared experts to capture common knowledge and reduce redundancy in routed experts. MoE’s success has also extended to the vision domain. LIMoE (mustafa2022multimodalcontrastivelearninglimoe) replaces dense MLP layers with MoE layers in CLIP, improving zero-shot image classification. AdaMV-MoE (liu2024adamolefinetuninglargelanguage) introduces an adaptive MoE framework for multi-task learning. Furthermore, CuMo (li2024cumoscalingmultimodalllm) integrates Co-upcycled Top-K sparse-gating MoE blocks into the visual encoder, MLP connector, and language model, enhancing VLLM inference performance.
@@
To address a critical gap where current VLLMs still struggle to solve vision-extensive multimodal tasks, we present our framework, MoCHA, in this section. Building upon the mainstream LLaVA architecture, MoCHA moves beyond the original CLIP encoder by integrating a diverse set of vision encoders, spanning different architectures and pre-training objectives. This enables the model to effectively capture and process visual information across varying levels of granularity. As illustrated in Figure 1, our framework consists of multiple vision encoders (i.e., OpenAI CLIP ViT-L/14@336, SigLIP ViT-L/16@384, OpenCLIP ConvNeXt-XXL@1024 and DINOv2 ViT-L/14@336) connected by a mixture of experts module and hierarchical group attention.
@@
Although language-supervised models outperform self-supervised and other models across all benchmark categories, Cambrian-1 (tong2024cambrian1fullyopenvisioncentric) highlights that well-trained self-supervised models like DINOv2 are capable of achieving competitive performance in vision-centric tasks. In particular, high-resolution models excel in chart-related and other vision-centric benchmarks while also demonstrating robust performance across both general VQA and knowledge-based VQA. While ViTs remain the dominant architecture, ConvNet-based models such as OpenCLIP ConvNeXt are well-suited for high-resolution image processing (vishniakov2024convnetvstransformersupervised) delivering outstanding results on OCR, chart interpretation, and other vision-centric benchmarks.
@@
Previous mainstream approaches (shazeer2017outrageouslylargeneuralnetworks) replace dense MLP blocks with sparsely-gated mixture of experts blocks. As an example, consider the SigLIP vision encoder embedding. Given input $e_{s}\in\mathbb{R}^{s\times D_{s}}$ and an MLP block, the hidden representation is computed as:
@@
While channel concatenation (lin2023sphinxjointmixingweights) achieves performance comparable to sequence append (liu2024prismervisionlanguagemodelmultitask; kar2024bravebroadeningvisualencoding; fan2024mousipolyvisualexpertvisionlanguagemodels), it requires strict spatial alignment of encoder outputs, rendering it less compatible with encoders of differing resolutions and architectures. Sequence append along the token dimension obviates token count alignment and removes the requirement for interpolation, resampling, or padding. As a result, encoders with different resolutions, patch sizes, or feature map dimensions can contribute arbitrary numbers of tokens. In this study, we adopt sequence append, concatenating features from different encoders directly along the token dimension:
@@
To maintain a load balance among experts in each MoEC block, we adopt auxiliary losses based on the language modeling cross-entropy loss. This auxiliary loss consists of a load balancing loss (wang2024auxiliarylossfreeloadbalancingstrategy) and a router z-loss (zoph2022stmoedesigningstabletransferable). Hence, the total loss is:
@@
 Qwen-VL-Chat  (bai2023qwenvlversatilevisionlanguagemodel)  Qwen-7B    57.5 68.2 61.5 - - 1487.5 60.6 -
 InstructBLIP  (dai2023instructblipgeneralpurposevisionlanguagemodels)  Vicuna-7B    49.2 60.5 50.1 26.2 - - 36.0 -
 LLaVA-LLaMA3 (contributorsxtuner)  LLaMA3-8B-IT    62.6 72.9 59.0 - 86.4 1469 72.3 -
 Mini-Gemini (li2024minigeminiminingpotentialmultimodality)  Vicuna-7B    - 65.2 - 40.8 - 1523 69.3 31.4
 SPHINX-Intern2 (liu2025sphinxxscalingdataparameters)  InternLM2-7B    56.2 70.4 58.1 36.5 86.9 1260.4 57.9 35.5
 LLaVA-NeXT (liu2024llavanext)  Vicuna-7B    64.2 70.1 64.9 43.9 86.5 1519 67.4 34.6
 LLaVA-NeXT (liu2024llavanext)  Mistral-7B    64.8 72.8 65.7 47.3 86.7 1498 68.7 37.7
 LLaVA-v1.5 (liu2023visual)  Vicuna-7B    62.0 66.8 58.2 30.5 85.9 1510.7 64.3 -
 VILA (lin2024vilapretrainingvisuallanguage)  Vicuna-7B    62.3 68.2 64.4 34.9 85.5 1533 68.9 -
 CuMo (li2024cumoscalingmultimodalllm)  Mistral-7B    64.9 73.9 67.0 51.0 86.7 1548.6 73.0 35.1
 InstructBLIP (dai2023instructblipgeneralpurposevisionlanguagemodels)  Vicuna-13B    49.5 63.1 50.7 25.6 78.9 1212.8 - -
 LLaVA-v1.5 (liu2023visual)  Vicuna-13B    63.3 71.6 61.3 35.4 85.9 1531.3 67.7 27.6
 Mini-Gemini (li2024minigeminiminingpotentialmultimodality)  Vicuna-13B    - 65.9 - 46.0 - 1565 68.5 37.0
 InternVL-Chat (chen2024internvl)  Vicuna-13B    66.6 - 61.5 - 87.6 1586.4 - -
 LLaMA-VID (li2024llama)  Vicuna-13B    65.0 70.0 - - 86.0 1542.3 66.6 -
 SPHINX-Plus (liu2025sphinxxscalingdataparameters)  LLaMA2-13B    - 74.2 65.7 47.9 89.1 1457.7 71.0 36.8
 MobileVLM (chu2023mobilevlmfaststrong)  MobileLLaMA-2.7B    59.0 61.0 47.5 - 84.9 1288.9 59.6 -
 TinyGPT-V (yuan2024tinygptvefficientmultimodallarge)  Phi2-2.7B    33.6 - - - - - - -
 LLaVA-Phi (zhu2024llava)  Phi2-2.7B    - 68.4 48.6 28.9 85.0 1335.1 59.8 -
 MoE-LLaVA-2.7B×4-Top2 (lin2024moellavamixtureexpertslarge)  Phi2-2.7B    61.4 68.5 51.4 34.3 86.3 1423 65.2 -
@@
We focus on academic VQA datasets, including GQA (hudson2019gqa), Science-QA (lu2022learn), and TextVQA (singh2019towards), as well as instruction-following VLLM benchmarks, such as POPE (li2023evaluatingobjecthallucinationlarge), MME (fu2024mmecomprehensiveevaluationbenchmark), MMBench (liu2024mmbenchmultimodalmodelallaround), and MM-Vet (yu2024mmvetevaluatinglargemultimodal). Furthermore, we evaluate the challenging MathVista (lu2024mathvistaevaluatingmathematicalreasoning) dataset to assess VLLM visual reasoning capabilities.
@@
As we scale up the LLM backbone, MoCHA (Vicuna-7B) demonstrates stronger visual reasoning capabilities than MoCHA (Phi-2 2.7B). Notably, MoCHA (Vicuna-7B) also surpasses models with even larger LLMs across multiple visual benchmarks. For instance, compared to LLaVA-v1.5 (Vicuna-13B), it achieves improvements of 2.99%, 4.42%, and 4.08% on SQA IMG, TextVQA, and POPE, respectively. Moreover, in contrast to CuMo—which applies a global mixture of experts (MoE) across the LLM, vision encoder, and connector—MoCHA (Vicuna-7B) outperforms CuMo (Mistral-7B) on several key vision benchmarks. These results provide strong evidence that, compared to the conventional single-vision-encoder paradigm, MoCHA’s dynamic MoE design and customized attention mechanisms offer a more effective approach for multimodal learning (ML). We present more results regarding other visual capabilities such as multi-object recognition in Appendix D.
@@
To validate the efficiency use of our MoECs, we further present training GFLOPs, parameters, and average inference time of models in Table 2. We make sure that all numbers are obtained under identical configurations to ensure the validity of the comparison. From the table, MoCHA achieves an inference speed of 0.57s with only 4.97B parameters and 12014.64 GFLOPs. This not only significantly reduces both the parameter size and computational cost, but also outperforms LLaVA-v1.5 (Vicuna-13B) and InternVL-Chat (Vicuna-13B) in inference efficiency. Furthermore, it approaches the inference speed of MoE-LLaVA, despite using even fewer parameters. These notable advantages in both computational and parameter efficiency highlight the strong potential of scaling up MoCHA as a compelling alternative to existing vision encoder paradigms in ML.
@@
Ablation Studies
Effect of Channel v.s. Sequential Concatenation.
@@
SigLIP+CLIP+ConvNeXt    59.10 66.58 87.06 60.83
@@
To evaluate MoEC’s capability in efficiently integrating heterogeneous and multi-dimensional features, we initiate this ablation study by replacing each MLP connector with an upcycled MoEC, as shown in Table 4. The results in Table 3 indicate that with the plain MLP as the connector, the model performance with a single SigLIP remains weak. While the addition of ConvNeXt further improves performance, the full integration of CLIP yields the largest gain among all settings. On the other hand, in Table 4, the combination of SigLIP, DINOv2, ConvNeXt, and CLIP under our MoEC results in notable performance improvements of 1.49%, 1.91%, 0.54%, and 1.52% on the GQA, SQA IMG, POPE, and MMBench benchmarks, respectively, compared to results using vanilla MLP in Table 3. These findings show that MoECs effectively handle diverse visual features, including object recognition, scene understanding, attribute identification, and spatial reasoning. Even as the number of vision encoders grows, MoECs incur minimal parameter overhead while keeping inference time and computational cost on par with a standard MLP. Further quantitative results are provided in Appendix C.
Method (MoEC + HGA)    GQA